# SAQR-QC: A Logic for Scalable but Approximate Quantitative Reasoning about Quantum Circuits

### v1

@@ Abstract.
Reasoning about quantum programs remains a fundamental challenge, regardless of the programming model or computational paradigm. Despite extensive research, existing verification techniques are insufficient—even for quantum circuits, a deliberately restricted model that lacks classical control, but still underpins many current quantum algorithms. Many existing formal methods require exponential time and space to represent and manipulate (representations of) assertions and judgments, making them impractical for quantum circuits with many qubits. This paper presents a logic for reasoning in such settings, called SAQR-QC. The logic supports Scalable but Approximate Quantitative Reasoning about Quantum Circuits, whence the name. SAQR-QC has three characteristics: (i) some (deliberate) loss of precision is built into it; (ii) it has a mechanism to help the accumulated loss of precision during a sequence of reasoning steps remain small; and (iii) most importantly, to make reasoning scalable, all reasoning steps are local—i.e., they each involve just a small number of qubits. We demonstrate the effectiveness of SAQR-QC via two case studies: the verification of GHZ circuits involving non-Clifford gates, and the analysis of quantum phase estimation—a core subroutine in Shor’s factoring algorithm.
@@ 1. Introduction
Quantum computing, as a new computational paradigm, exploits quantum parallelism and interference to outperform classical methods on specific tasks. For instance, Shor’s algorithm (Shor94, ) efficiently factors large integers using the Quantum Fourier Transform (QFT). Kitaev’s phase estimation (Kitaev1995, ) serves as a foundation for quantum simulation (Aspuru-Guzik2005, ) and optimization algorithms, such as QAOA (FarhiGoldstoneGutmann14, ). Interestingly, key components that provide the advantages of quantum computation over classical computation are obtained (1) without the use of classical control-flow constructs such as while loops, and (2) with a fixed number of qubits, the quantum analogue of bits, throughout the computation.
@@
Despite quantum computing being a much different computational paradigm, one is still interested in showing that programs behave as desired. Significant effort has been put into developing methods to verify quantum programs using classical computers (BHY19, ; ZYY19, ; zhou2021quantum, ; 10.1145/3453483.3454029, ; Unruh, ; YP21, ; DBLP:journals/pacmpl/0001CLLTY23, ; 10.1145/3704873, ; 10.1145/3519939.3523431, ). Particular noteworthy is Ying’s Quantum Hoare Logic (QHL) (Ying11, ), which extends classical Hoare logic to reason about the partial correctness of quantum programs, using quantum predicates, semidefinite positive Hermitian operators, as preconditions and postconditions. The quantum Hoare triple, introduced in (DP06, ), is defined as
(1) $\{A\}P\{B\}\quad\mathrm{iff}\quad\textrm{for all}~{}\rho,{\mathrm{Tr}}(A\rho)\leq{\mathrm{Tr}}(B\llbracket P\rrbracket(\rho)),$
where $A$ and $B$ are quantum observables representing preconditions and postconditions (and ${\mathrm{Tr}}$ is the matrix-trace operation).
Despite this extensive body of research, quantum program verification remains inadequate, even for quantum circuits—a deliberately restricted model that lacks classical control in which many current quantum algorithms are formulated. Existing formal methods struggle to scale; for instance, although QHL provides a comprehensive framework for tracking quantum program behavior, including success probabilities, it lacks support for local reasoning and requires exponential time and space, making it impractical for large systems. More precisely, for a unitary program $P$, whose semantics corresponds to a unitary matrix $U_{P}$, QHL provides the rule
(2) $\{U_{P}^{{\dagger}}BU_{P}\}P\{B\}$
Although this rule is complete for unitary programs, it imposes a computationally intractable task on the verification process—namely, computing $U_{P}^{\dagger}BU_{P}$ for a general unitary $U_{P}$. Unfortunately, even representing a general $n$-qubit unitary matrix (or a general observable) requires $(2^{n})^{2}=4^{n}$ parameters.
This discussion raises the following question: {mdframed} Is there a theoretical foundation for scalable quantitative reasoning about quantum programs that use many qubits? Here, “scalable” means that the computational resources required for verification grow polynomially with the number of qubits.
This paper answers this question in the affirmative. It presents a logic for reasoning about programs expressed as quantum circuits, called SAQR-QC—for Scalable but Approximate Quantitative Reasoning about Quantum Circuits. SAQR-QC has three characteristics: (i) some (deliberate) loss of precision is built into it; (ii) it has a mechanism to help the accumulated loss of precision during a sequence of reasoning steps remain small; and (iii) most importantly, to make reasoning scalable, all reasoning steps are local—i.e., they each involve just a small number of qubits.
SAQR-QC applies to the circuit model of quantum programs, which is widely used in quantum computing. Most known quantum algorithms either fit within the quantum-circuit model or consist of some number of runs of a quantum circuit. We consider quantum circuits that act on $n$ qubits and are composed of a polynomial number of unitary gates of up to $k$ qubits (where $k$ is a constant independent of $n$).111 We place no restrictions on the gate set, except that each elementary gate operates on a constant number of qubits. While our examples use one- and two-qubit gates, our results extend to circuits composed of gates acting on three qubits (such as the Toffoli gate), four qubits, etc. With quantum circuits, the input state is typically initialized to some computational basis state, such as |$\ket{0}^{\otimes n}$.
Does this setting —-quantum circuits composed of a polynomial number of $k$-qubit gates acting on |$\ket{0}^{\otimes n}$ —-make the problem simple enough to be efficiently solved by a classical computer? Not at all. Even if we only care about the value of the output state’s first qubit, which can be represented by a $2\times 2$ density matrix, and even if we restrict our interest to just the first diagonal entry of that matrix, the problem remains hard. Determining whether this single number is smaller than $1/3$ or larger than $2/3$ is believed not to be solvable by any classical polynomial-time algorithm unless quantum computation offers no advantage over classical computation.
One of the inspirations for our work is Quantum Abstract Interpretation (QAI) (YP21, ), a framework proposed by Yu and Palsberg that extends classical abstract interpretation (DBLP:conf/popl/CousotC77, ) to reason about quantum circuits. A QAI judgment for a quantum circuit $C$ is formalized as follows:
@@
(3) $\models^{\rm QAI}\{\mathcal{P}\}C\{\mathcal{Q}\}\quad\mathrm{iff}\quad\textrm{for all}~{}\rho,\rho\vDash\mathcal{P}~{}\mathrm{implies}~{}\llbracket\mathbf{C}\rrbracket(\rho)\vDash\mathcal{Q},$
where $\mathcal{P},\mathcal{Q}$ are tuples of local projections used as abstract states. The above triple asserts that if the program’s input satisfies $\mathcal{P}$, the output will satisfy $\mathcal{Q}$, where satisfaction means that the output state lies within the space defined by the intersection of the spaces of the local projections in $\mathcal{Q}$.
QAI facilitates the analysis and verification of quantum programs by (a) mapping quantum states to an abstract domain that capture essential properties of the quantum state, and (b) providing a technique for applying abstract transformers in which all reasoning is local—i.e., each application of an abstract transformer involves just a small number of qubits. On the plus side, QAI enables polynomial-time analysis. However, on the minus side, QAI does not support the reasoning of quantitative. The local-reasoning method in QAI focuses on local projections, which provide qualitative insight but fail to capture success probabilities. However, most quantum algorithms succeed with probability less than 1, underscoring the need for a scalable framework for quantitative analysis of quantum programs—a challenge that remains unaddressed. Here, scalability requires verification algorithms that are significantly more efficient in time and space than brute-force simulation, which scales exponentially with the number of qubits.
QAI and QHL stand on opposite ends of a feature spectrum: whereas QAI is scalable but lacks quantitative reasoning, QHL offers quantitative reasoning, but is not scalable. SAQR-QC takes inspiration from both. Integrates QHL-like quantitative reasoning with the scalability of QAI to achieve scalable quantitative reasoning about quantum circuits with |$\ket{0}^{\otimes n}$ as input state.
In SAQR-QC, the idea behind the QHL-like component is to replace a single observable with a tuple of local observables, each represented by a semidefinite matrix that captures information about the corresponding reduced density matrix (NI11, ).
Notably, a tuple of reduced density matrices, each defined on a constant number of qubits, requires only a linear number of parameters to describe, contrasting sharply with the exponential complexity needed to represent the full quantum state. This locality enables tractable analysis and verification techniques, making it feasible to reason about large-scale quantum computations through their low-dimensional marginals. A judgment in SAQR-QC takes the following form:
@@
Here, $\mathscr{A}$ and $\mathscr{B}$ are tuples of local observables used to track the probability of success, where locality means that each observable acts non-trivially on only a constant number of qubits, independent of the total number of qubits. This locality ensures the potential for scalable reasoning in a manner similar to that of QAI. Likewise, $\mathcal{P}$ and $\mathcal{Q}$ are tuples of local projections, as defined in QAI. 222 The two-part structure of each assertion is analogous to how formulas in reasoning systems based on separation logic have two parts—a spatial formula and a pure logical formula) (DBLP:conf/fmco/BerdineCO05, ; DBLP:conf/tacas/DistefanoOY06, ). In both cases, the principle is to carry around two different kinds of constraints on the state. In SAQR-QC, the QAI-like component provides spatial information about the program state—i.e., the quantum state lies within a restricted subspace. The local observables of the QHL-like component then provide a bound on the inner product between the program output and a local observable. More precisely, the triple in Equation 4 asserts
@@
(5) $\textrm{for all}~{}\rho,\rho\vDash\mathcal{P}~{}\mathrm{implies}~{}(i)~{}\llbracket\mathbf{C}\rrbracket(\rho)\vDash\mathcal{Q},~{}~{}\mathrm{and}~{}~{}(ii)~{}{\mathrm{Tr}}(M_{\mathscr{A}}\rho)\leq{\mathrm{Tr}}(M_{\mathscr{B}}\llbracket\mathbf{C}\rrbracket(\rho)),$
@@
where $M_{\mathscr{A}}$ and $M_{\mathscr{B}}$ are derived from $\mathscr{A}$ and $\mathscr{B}$, with the precise definition provided in § 4.
SAQR-QC has built into it some (deliberate) loss of precision, but the use of mechanisms from both QHL and QAI causes the accumulated loss of precision during a sequence of reasoning steps to remain small. Both the deliberate-loss-of-precision aspect and the use of two interacting formalisms are inspired by abstract interpretation.
An important use-case that we envision for SAQR-QC is for forward reasoning—i.e., given the pre-state assertion $\{\mathscr{A}|\mathcal{P}\}$333 Even though we restrict attention to circuits with input |$\ket{0}^{\otimes n}$, stating the pre-state assertion $\{\mathscr{A}|\mathcal{P}\}$ is not necessarily trivial. For the $\mathscr{A}$ component, one needs to choose what local observables to track; For the $\mathcal{P}$ component, one needs to choose what local projections to track. and circuit $C$, find a suitable post-state assertion $\{\mathscr{B}|\mathcal{Q}\}$. Obviously, the post-state assertion obtained using SAQR-QC will not characterize the strongest post-condition of $C$ with respect to $\mathscr{A}|\mathcal{P}$. Instead, the goal is that the obtained post-condition will be strong enough to obtain useful information about the computation.
One important instance of such information is the computation’s success probability, which, in models like BQP,444 “BQP is the class of decision problems solvable by a quantum computer in polynomial time with bounded error,” just as “P is the class of decision problems solvable by a classical deterministic computer in polynomial time.” is typically determined by a diagonal entry of the output state’s first qubit.
It should be borne in mind that SAQR-QC is a logic, not an automated reasoning system. SAQR-QC is similar to other logics in that clever insights and eureka steps are sometimes needed to be able to establish that certain properties do, in fact, hold. One of our goals is to be able to establish success probabilities for quantum circuits, and doing so with SAQR-QC is not “cookbook.” For instance, the two case studies presented in § 5 and § 6 are both examples in which the specific choices of the predicates used are crucial to obtaining the sharp results detailed in those sections.
@@
The work described in the paper makes the following contributions:
  • We present SAQR-QC, which supports—and is named for—Scalable but Approximate Quantitative Reasoning about Quantum Circuits.
  • SAQR-QC takes inspiration from QHL: similar to QHL, it supports quantitative reasoning.
  • SAQR-QC takes inspiration from QAI: similar to QAI, some (deliberate) loss of precision is built into the reasoning rules of SAQR-QC. The advantage is the same as in QAI, namely that because all reasoning steps are local—i.e., they each involve just a small number of qubits—they can be performed in polynomial time, and thus the reasoning method is scalable.
  • SAQR-QC takes inspiration from abstract interpretation: in particular, the use of the QAI-like information to restrict the QHL-like information is a mechanism to help the accumulated loss of precision during a sequence of reasoning steps remain small, and hence similar to a reduced product (DBLP:conf/popl/CousotC79, , §10.1).
  • We demonstrate the effectiveness of SAQR-QC via two case studies:
    • The verification of GHZ circuits involving non-Clifford gates: We apply SAQR-QC to derive a judgment that gives a characterization of the output state that is precise in every aspect except for a phase factor.
    • The analysis of quantum phase estimation—a core subroutine in Shor’s factoring algorithm: We apply SAQR-QC to derive a judgment showing that, for a given constant $k$, the algorithm provides the best estimation of the last $k$ bits of the phase with probability at least $\frac{4}{\pi^{2}}$. To obtain this result, we first develop a lossless local-reasoning method for the quantum Fourier transform using QAI. In particular, the concretization of the abstract output state exactly matches the true output state for computational-basis inputs. To the best of our knowledge, no prior approach has shown that within polynomial time it is possible to obtain such results about either the quantum Fourier transform or quantum phase estimation.
@@
We present background material about quantum computing in § 2. § 3 summarizes the basics of QAI (YP21, ). § 4 defines local observables as predicates, and formalizes the judgments of SAQR-QC, which integrate local observables and QAI. We employ SAQR-QC to reason about the general GHZ circuit (§ 5) and quantum phase estimation (§ 6). § 7 discusses related work. § 8 concludes. Some proofs and derivations are available in the paper’s appendices, which are submitted as Supplementary Material.
@@ 2. Background & Notation
To make the paper self-contained, this section provides background material, and discusses notation for—and properties of—quantum computing. The material in this section is similar to what can be found in published books and papers, e.g., (NI11, ; YP21, ; Ying11, ). Readers already familiar with quantum computing may wish to proceed directly to § 3.
@@ 2.1. Preliminaries
We use the notation $[n]$ to denote the set $\{1,\ldots,n\}$, $\setminus$ to denote set difference, and $|s|$ to denote the cardinality of a set $s$. We assume familiarity with Dirac notation and standard linear algebra concepts, including Hilbert space, tensor products, orthonormal bases, and inner and outer products.
Linear operators on $d$-dimensional complex vector spaces are represented as $d\times d$ matrices over $\mathbb{C}$, denoted $\mathbb{C}^{d\times d}$. The identity operator is denoted by $I$. For an operator $A$, the conjugate transpose is defined as $A^{\dagger}=(A^{T})^{*}$, where $A^{T}$ is the transpose and $(\cdot)^{*}$ denotes complex conjugation. An operator $A$ is Hermitian if $A=A^{\dagger}$, and positive semi-definite if all its eigenvalues are non-negative. The trace of a matrix $A$, denoted ${\mathrm{Tr}}(A)$, is the sum of its diagonal entries: ${\mathrm{Tr}}(A)=\sum_{i}A_{ii}$.
The Löwner order, defined on Hermitian matrices in quantum mechanics and convex analysis, establishes a relationship $A\leq B$ indicating that $B-A$ is positive semidefinite. The Löwner order is vital for comparing states and operators.

### v3

# A Logic for Approximate Quantitative Reasoning about Quantum Circuits

@@ Abstract.
Reasoning about quantum programs remains a fundamental challenge, regardless of the programming model or computational paradigm. Despite extensive research, existing verification techniques are insufficient—even for quantum circuits, a deliberately restricted model that lacks classical control, but still underpins many current quantum algorithms. Many existing formal methods require exponential time and space to represent and manipulate (representations of) assertions and judgments, making them impractical for quantum circuits with many qubits. This paper presents a logic for reasoning in such settings, called SAQR-QC. The logic supports Scalable but Approximate Quantitative Reasoning about Quantum Circuits, whence the name. SAQR-QC has three characteristics: (i) some (deliberate) loss of precision is built into it; (ii) it has a mechanism to help the accumulated loss of precision during a sequence of reasoning steps remain small; and (iii) most importantly, to make reasoning scalable, every reasoning step is local—i.e., it involves just a small number of qubits. We demonstrate the effectiveness of SAQR-QC via two case studies: the verification of GHZ circuits involving non-Clifford gates, and the analysis of quantum phase estimation—a core subroutine in Shor’s factoring algorithm.
@@ 1. Introduction
Quantum computing leverages superposition and interference to achieve computational advantages over classical methods for specific tasks. For instance, Shor’s algorithm (Shor94, ) efficiently factors large integers using the Quantum Fourier Transform (QFT) and quantum phase estimation (Kitaev1995, ). The quantum-phase-estimation technique is also a foundational component in quantum-simulation algorithms (Aspuru-Guzik2005, ) and in quantum algorithms for solving linear algebra problems (HHL08, ).
@@
Despite the different computational paradigm, ensuring that quantum programs behave as intended remains a critical challenge. This challenge has spurred significant research into verifying quantum programs using classical computers (BHY19, ; ZYY19, ; zhou2021quantum, ; 10.1145/3453483.3454029, ; Unruh, ; YP21, ; 10.1145/3704873, ; 10.1145/3519939.3523431, ; CACM:CCLLTY25, ). A landmark achievement is Ying’s Quantum Hoare Logic (QHL) (Ying11, ), which extends classical Hoare logic to reason about the partial correctness of quantum programs. QHL uses quantum predicates—semidefinite positive Hermitian operators—as pre- and postconditions, with the quantum Hoare triple (DP06, ) defined as:
(1) $\{A\}P\{B\}\quad\mathrm{iff}\quad\textrm{for all}~\rho,{\mathrm{Tr}}(A\rho)\leq{\mathrm{Tr}}(B\llbracket P\rrbracket(\rho)).$
Despite this progress, scalable verification remains elusive, even for quantum circuits—a restricted model without classical control that encapsulates many current algorithms. Existing methods struggle with exponential complexity; for instance, while QHL is complete for unitary programs, its application requires computing $U_{P}^{\dagger}BU_{P}$ for a general $n$-qubit unitary $U_{P}$, a task requiring time and space exponential in $n$. This intractability motivates our central question: Can we establish a theoretical foundation for scalable quantitative reasoning about quantum programs that use many qubits? Here, “scalable” means that the size of a proof—the matrices involved and the logical derivation itself—grows polynomially with the number of qubits.
This paper answers this question affirmatively by introducing SAQR-QC (Scalable but Approximate Quantitative Reasoning for Quantum Circuits), a logic for reasoning about programs expressed as quantum circuits. SAQR-QC is designed with three principles: (i) it embraces a deliberate, controlled loss of precision to achieve scalability; (ii) it incorporates mechanisms to keep the accumulated imprecision small across reasoning steps; (iii) all reasoning is local, meaning each step involves only a constant number of qubits, independent of the total system size.
SAQR-QC is inspired by two prior approaches that sit at opposite ends of a spectrum. At one end, QHL—see Equation˜1—offers fully precise quantitative reasoning, but is computationally intractable for large systems. At the other end, Quantum Abstract Interpretation (QAI) (YP21, ) provides a scalable, qualitative framework. A QAI judgment,
@@
$\models^{\rm QAI}\{\mathcal{P}\}C\{\mathcal{Q}\}\quad\mathrm{iff}\quad{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\mathrm{forall}\,\rho,\,\rho\vDash\mathcal{P}~\text{implies}~\llbracket\mathbf{C}\rrbracket(\rho)\vDash\mathcal{Q},}}$
uses tuples of local projections $\mathcal{P}$ and $\mathcal{Q}$ as abstract states. QAI enables efficient, local reasoning but lacks the ability to reason about quantitative properties like success probabilities.
SAQR-QC integrates ideas from both to achieve scalable quantitative reasoning. It is important to clarify that SAQR-QC is a logic for manual proof construction, analogous to how Hoare Logic, Separation Logic, and Linear Logic were first developed as conceptual frameworks to facilitate human-driven reasoning. The quest for automation, while a vital future direction, is not the aim of this paper.
A judgment in SAQR-QC has a two-part structure, inspired by the spatial and pure assertions of separation logic (DBLP:conf/fmco/BerdineCO05, ; DBLP:conf/tacas/DistefanoOY06, ):
@@
Here, $\mathcal{P}$ and $\mathcal{Q}$ are QAI-style tuples of local projections, capturing qualitative spatial constraints on the state. $\mathscr{A}$ and $\mathscr{B}$ are tuples of local observables, each acting non-trivially on only a constant number of qubits. These observables track quantitative information, such as success probabilities, in a manner reminiscent of QHL, but restricted to local components. The triple (2) asserts that for any input state $\rho$ satisfying $\mathcal{P}$, the output state $\llbracket C\rrbracket(\rho)$ satisfies $\mathcal{Q}$, and a certain quantitative relationship involving $\mathscr{A}$ and $\mathscr{B}$ holds. More precisely, a SAQR-QC judgment means
@@
(3) $\models^{\rm\textsc{SAQR-QC}}\{\mathscr{A}|\mathcal{P}\}C\{\mathscr{B}|\mathcal{Q}\}\quad\mathrm{iff}\quad\textrm{for all}~\rho,\,\rho\vDash\mathcal{P}~\textrm{implies}~~\begin{aligned} &\textrm{(i)}~\llbracket\mathbf{C}\rrbracket(\rho)\vDash\mathcal{Q},~~\mathrm{and}\\ &\textrm{(ii)}~{\mathrm{Tr}}(M_{\mathscr{A}}\rho)\leq{\mathrm{Tr}}(M_{\mathscr{B}}\llbracket\mathbf{C}\rrbracket(\rho)),\end{aligned}$
@@
where $M_{\mathscr{A}}$ and $M_{\mathscr{B}}$ are derived from $\mathscr{A}$ and $\mathscr{B}$, with the precise definition provided in §˜3.
This two-part structure is not merely stylistic; it is essential under the locality constraint. Without locality, the two components could be combined into a single QHL triple (ZYY19, , Theorems 3.2 and 3.3). However, with the restriction to local projections and observables, this combination is generally impossible, justifying the distinct roles of the QAI-like and QHL-like components. Their interaction—inspired by the reduced product in abstract interpretation (DBLP:conf/popl/CousotC79, , §10.1), as well as the use of abstract interpretation in logic-based tools (Crab, )—helps to control the accumulation of imprecision.
A primary use-case for SAQR-QC is forward reasoning: given a pre-state assertion $\{\mathscr{A}|\mathcal{P}\}$ and a circuit $C$, derive a post-state assertion $\{\mathscr{B}|\mathcal{Q}\}$. The derived post-condition might not be the strongest possible, but aims to be strong enough to extract useful quantitative information, such as the success probability of a computation in the BQP model. [^1] BQP is the class of decision problems solvable by a quantum computer in polynomial time with bounded error, analogous to how P is defined for classical deterministic computers.
As with any manual logic, constructing a SAQR-QC proof requires insight, particularly in choosing the appropriate local projections and observables. The logic provides the framework, but the user must supply the clever insights—the “eureka” steps—to define the predicates that make the verification possible. The case studies in §˜4 and §˜5 exemplify this principle, demonstrating how specific, carefully chosen predicates yield sharp results for non-trivial quantum circuits.
@@
This paper makes the following contributions:
  • We present SAQR-QC, a logic for Scalable but Approximate Quantitative Reasoning about Quantum Circuits. SAQR-QC integrates quantitative, QHL-like reasoning with the scalable, local reasoning of QAI.
  • We provide a formal foundation for SAQR-QC, defining its assertions and proof rules. The design ensures that all reasoning steps are local, leading to proofs whose size is polynomial in the number of qubits and gates.
  • We demonstrate the logic’s utility and expressiveness through two detailed case studies:
    • Verification of GHZ circuits involving non-Clifford gates: We derive a judgment that characterizes the output state precisely in every aspect except a relative phase factor.
    • Analysis of Quantum Phase Estimation (QPE): We derive a judgment showing that for any constant $k$, the QPE algorithm provides the best estimate of the last $k$ bits of the phase, with probability at least $4/\pi^{2}$. A key step is a novel, lossless local-reasoning method for the Quantum Fourier Transform using QAI. To the best of our knowledge, no prior approach has achieved such a result with a proof that scales polynomially with the system size.
  • We clarify the relationship between SAQR-QC and prior work, particularly Zhou et al. (ZYY19, ), and discuss the principles for constructing effective predicates for use in our framework.
In summary, this work takes a foundational step towards scalable reasoning about quantum computations by defining a logic that is both quantitative and tractable, and by demonstrating its power on representative quantum algorithms.
@@
§˜2 presents background material about quantum computing, as well as the basics of QAI. §˜3 defines local observables as predicates, and formalizes the judgments of SAQR-QC, which integrate local observables and QAI. We employ SAQR-QC to reason about the general GHZ circuit (§˜4) and quantum phase estimation (§˜5). §˜6 discusses related work. §˜7 concludes. Some proofs and derivations are available in the paper’s appendices, which are submitted as Supplementary Material.
@@ 2. Background & Notation
To make the paper self-contained, this section reviews basic concepts, notation, and properties of quantum computing, following standard references such as (NI11, ; YP21, ; Ying11, ), as well as the basics of QAI. Readers familiar with these topics may wish to skip to §˜3.
@@ 2.1. Preliminaries
[…]
Linear operators on $d$-dimensional complex vector spaces are represented by $d\times d$ matrices $\mathbb{C}^{d\times d}$. The identity is $I$, and the conjugate transpose of $A$ is $A^{\dagger}=(A^{T})^{*}$. An operator is Hermitian if $A=A^{\dagger}$ and positive semi-definite if all eigenvalues are non-negative. The trace, ${\mathrm{Tr}}(A)$, is the sum of the diagonal entries: ${\mathrm{Tr}}(A)=\sum_{i}A_{ii}$.
The Löwner order on Hermitian matrices, $A\leq B$ if BAB-A is positive semidefinite, is fundamental for comparing states and operators in quantum mechanics.